Depreciation Methods

Gruman Company purchased a machine for $198,000 on January 2, 2019. It made the following estimates:

Service life 5 years or 10,000 hours
Production 180,000 units
Residual value $ 18,000

In 2019, Gruman uses the machine for 2,000 hours and produces 45,000 units. In 2020, Gruman uses the machine for 1,400 hours and produces 34,000 units. If required, round your final answers to the nearest dollar.

Required:

  1. Compute the depreciation for 2019 and 2020 under each of the following methods:
  1. Straight-line method
    2019 $
    2020 $
  2. Sum-of-the-years'-digits method
    2019 $
    2020 $
  3. Double-declining-balance method
    2019 $
    2020 $
  4. Activity method based on hours worked
    2019 $
    2020 $
  5. Activity method based on units of output
    2019 $
    2020 $
  1. For each method, what is the book value of the machine at the end of 2019? At the end of 2020?
  1. Straight-line method
    2019 $
    2020 $
  2. Sum-of-the-years'-digits method
    2019 $
    2020 $
  3. Double-declining-balance method
    2019 $
    2020 $
  4. Activity method based on hours worked
    2019 $
    2020 $
  5. Activity method based on units of output
    2019 $
    2020 $
  1. If Gruman used a service life of 8 years or 15,000 hours and a residual value of $9,000 , what would be the effect on the following under the straight-line, sum-of-the-years'-digits, and double-declining-balance depreciation methods?

Depreciation expense

  1. Straight-line method
    2019 $
    2020 $
  2. Sum-of-the-years'-digits method
    2019 $
    2020 $
  3. Double-declining-balance method
    2019 $
    2020 $

Book value

  1. Straight-line method
    2019 $
    2020 $
  2. Sum-of-the-years'-digits method
    2019 $
    2020 $
  3. Double-declining-balance method
    2019 $
    2020 $
Answer #1

A. STRAIGHT LINE DEPRECIATION = COST- RESIDUAL VALUE/USEFUL LIFE OF ASSETS

2019= $198,000-$18,000/5 YEARS

= $36,000

2020= $36,000

B. SUM OF YEARS DIGIT METHOD

DEPRECATION EXPENSE = REMAINING USEFUL LIFE OF ASSETS*DEPRECIABLE COST/SUM OF YEARS

2019= 5*$ 180,000/15

= $60,000

2020 = 4* $ 180,000/15

= $ 48,000

C.DOUBLE DECLINING BALANCE METHOD

RATE= $198,000/5 = 39,600

= 39,600/198,000= 20%*2=40%

2019= $198,000*40%= $79,200

2020=( $198,000-79,200)*40%= $47,520

D. ACTIVITY BASED ON HOURS = DEPRECIABLE COST * HOURS WOEKED/TOTAL SERVICE LIFE

2019= $180,000*2,000/10,000

= $ 36,000

2020= $ 180,000*1,400/10,000

= $ 25,200

E. ACTIVITY METHOD BASED ON UNITS = DEPRECIABLE COST * PRDOUCTION IN YEAR/TOTAL PRODUCTION

2019= $ 180,000* 45,000UNITS/180,000 UNITS

= $ 45,000

2020= $ 180,000* 34,000 UNITS/180,000 UNITS

= $34,000

BOOK VALUE OF MACHINE

STRAIGHT LINE METHOD

YEAR BOOK V AT BEGINNING DEPRECIATION ACCUMULATED BOOK VALUE ENDING
2019 $ 198,000 36,000 36,000 162,000
2020 $162,000 36,000 72,000 126,000

B. SUM OF YEARS DIGIT METHOD

YEAR BOOK V AT BEGINNING DEPRECIATION ACCUMULATED BOOK VALUE ENDING
2019 $ 198,000 60,000 60,000 138,000
2020 $138,000 48,000 108,000 90,000

C. DOUBLE- DECLINING BALANCE METHOD

YEAR BOOK V AT BEGINNING DEPRECIATION ACCUMULATED BOOK VALUE ENDING
2019 $ 198,000 79,200 79,200 118,800
2020 $118,800 47,520 126,720 70,480

D. ACTIVITY METHOD BASED ON HOURS

YEAR BOOK V AT BEGINNING DEPRECIATION ACCUMULATED BOOK VALUE ENDING
2019 $ 198,000 36,000 36,000 162,000
2020 $162,000 25,200 61,200 136,800

E. ACTIVITY METHOD BASED ON UNITS

YEAR BOOK V AT BEGINNING DEPRECIATION ACCUMULATED BOOK VALUE ENDING
2019 $ 198,000 45,000 45,000 153,000
2020 $153,000 34,000 79,000 119,000

1. IF GRUMAN SERVICE LIFE.......

A. STRAIGHT LINE DEPRECIATION = COST- RESIDUAL VALUE/USEFUL LIFE OF ASSETS

2019= $198,000-$9,000/8 YEARS

= $23,625

2020= $23,625

B. SUM OF YEARS DIGIT METHOD

DEPRECATION EXPENSE = REMAINING USEFUL LIFE OF ASSETS*DEPRECIABLE COST/SUM OF YEARS

2019= 8*$ 189,000/36

= $ 42,000

2020 = 7* $ 189,000/36

= $ 36,750

C.DOUBLE DECLINING BALANCE METHOD

RATE= $198,000/8 = 24,750

= 24,750/198,000= 12.5%*2=25%

2019= $198,000*25%= $ 49,500

2020=( $198,000-49,500)*25%= $37,125

BOOK VALUE OF MACHINE

STRAIGHT LINE METHOD

YEAR BOOK V AT BEGINNING DEPRECIATION ACCUMULATED BOOK VALUE ENDING
2019 $ 198,000 23,625 23,625 174,375
2020 $174,375 23,625 47,250 127,125

B. SUM OF YEARS DIGIT METHOD

YEAR BOOK V AT BEGINNING DEPRECIATION ACCUMULATED BOOK VALUE ENDING
2019 $ 198,000 42,000 42,000 156,000
2020 $156,000 36,750 78,750 119,250

C. DOUBLE- DECLINING BALANCE METHOD

YEAR BOOK V AT BEGINNING DEPRECIATION ACCUMULATED BOOK VALUE ENDING
2019 $ 198,000 49,500 49,500 148,500
2020 $148,500 37,125 86,625 111,375